OpenStudy (anonymous):

Find the modulus of z if z=2e^(i*theta). Please help? Thanks so much!

OpenStudy (anonymous):

ie, |z| if\[z=2e^{i \theta}\]

OpenStudy (dumbcow):

\[\large 2e^{i \theta}=2(cos(\theta)+i \sin(\theta))\] \[|z| = 2\sqrt{\cos^{2} \theta + \sin^{2} \theta} = 2\]

OpenStudy (anonymous):

Thank you so much!